# Xizhou Zhu1,2*,  Xue Yang2*,  Zhaokai Wang3,2*,  Hao Li4,2 Wenhan Dou1,5,  Junqi Ge1,5,  Lewei Lu5,  Yu Qiao2,  Jifeng Dai1,2🖂

### v1

@@ Abstract
00footnotetext: Equal contribution. This work was completed by Zhaokai Wang and Hao Li during their internship at Shanghai Artificial Intelligence Laboratory. 🖂Corresponding author: Jifeng Dai <daijifeng@tsinghua.edu.cn>.
@@ 4.1 Implementation Details
For comparison with Base-size models, we use pre-trained ViT-T/S/B as the branches to construct three-branch PIIP network, namely PIIP-TSB. Similarly, ViT-S/B/L are used to construct PIIP-SBL to match the computation of Large-size models. We also construct four-branch PIIP-TSBL with ViT-T/S/B/L. We set the number of interactions (each with 2 interaction units as shown in Fig. 2) $N$ to 12, i.e. after every layer for ViT-T/S/B or after every two layers for ViT-L. We construct multiple variants of three-branch and four-branch models with different resolution configurations.
@@ 4.2 Object Detection and Instance Segmentation
Table 2: Object detection and instance segmentation performance on COCO val2017. ‘MS’ means multi-scale training. Large-size models use ViT weights trained on ImageNet-21K.
@@
DINO 3$\times$ + MS schedule
PIIP-SBL (ours) 57.9 76.9 63.3 - - -
@@
Effectiveness of Parameter-Inverted Image Pyramid. To demonstrate the performance and computational advantages of the Parameter-Inverted Image Pyramid (PIIP) Networks, we perform validation on two baseline models ViTDet-B and ViTDet-L [21] in Table 1. Taking the three-branch structure as an example, while maintaining similar performance with ViTDet-B, our PIIP-TSB reduces the computational cost by 47.5% (243G vs. 463G) and 38.0% (287G vs. 463G) in object detection and instance segmentation tasks respectively. Similarly, compared with ViTDet-L, our PIIP-SBL reduces the computational cost by about 52.9% (727G vs. 1,542G) and 35.0% (1,002G vs. 1,542G) in the above two tasks respectively. On the other hand, with similar computational cost as the baseline, PIIP-TSB and PIIP-SBL improve the object detection performance by 2.8% and 2.6%, respectively, and instance segmentation by 1.5% and 1.2%, compared to ViTDet-B and ViTDet-L. To better illustrate the above conclusion, we depict the trend between the computational cost and performance of different PIIP model combinations by adjusting the input resolution, as shown in Fig. 4. Furthermore, when we use the four-branch structures, the curve in the figure is slightly better than that of the three-branch structure.
@@
Results with Base-size and Large-size models. As shown in Table 2, combined with Mask R-CNN, PIIP achieves higher performance than ViT-Adapter by a considerable margin, about 0.9% and 1.2% on $\rm AP^{b}$. With a more powerful detector Cascade R-CNN and stronger training schedule (3$\times$ + MS), PIIP-TSB and PIIP-SBL achieve competitive performance of 53.1% and 54.5% $\rm AP^{b}$, respectively. Finally, we achieve 57.9% $\rm AP^{b}$ with the DINO [57] detector.
Results with InternViT-6B. We further examine PIIP on an extremely large vision foundation model InternViT-6B [8], and finally achieve 55.7% $\rm AP^{b}$ by using Mask R-CNN 1$\times$ training schedule. In addition, as can be seen from Table 3, our PIIP can save nearly 43% of the computation and achieve better performance than single-branch InternViT-6B by 1.9% on $\rm AP^{b}$ and 0.9% on $\rm AP^{m}$.
Results with different pre-training methods. We initialize PIIP-SBL with pre-trained ViT S/B/L weights from AugReg [40], Uni-Perceiver [62], MAE [16], DeiT III [43], DINOv2 [35] and BEiTv2 [36]. As shown in Table 6, the BEiTv2-initialized model achieves the best performance.
@@ 4.3 Semantic Segmentation
Results with Base-size and Large-size models. In Table 6, PIIP can achieve better performance with fewer computations. Moreover, PIIP-TSB in Table 6 attains 51.6% mIoU with UperNet, exceeding InternImage-B [46] by 1.4%. Similarly, PIIP-SBL yields the results of 54.3% mIoU, which is outstanding compared to counterparts like ConvNeXt-XL [29] and InternImage-L [46].
@@
Table 8: Ablation on Branch Merging with PIIP-TSB under resolution 1568/896/672.
@@ 4.4 Image Classification
Results. As shown in Table 8, when compared with the DeiT baseline, our PIIP-SBL reduces the computational cost by 36.7% (39.0G vs. 61.6G) while maintaining the performance. When using a similar computational cost as the baseline models, PIIP-TSB and PIIP-SBL improve the top-1 accuracy by 0.3% and 0.7%, respectively.
Table 9: Ablation on image pyramid and parameter-inverted design. ‘PI’ and ‘IP’, ‘Inter.’ represent parameter-inverted, image pyramid and interactions. ‘MS’ means multi-scale training.
@@ 4.5 Ablation Study
Superiority of parameter-inverted image networks. We evaluate the effectiveness of the image pyramid and parameter-inverted design by comparing our method with other methods, e.g. designs in Fig. 1. First of all, a single-branch with multi-scale training is the simplest image pyramid practice, as shown in Table 9. Compared with the baseline model, its performance improvement is limited (44.8% vs. 43.8%). Secondly, we conduct experiments by controlling the scale of the branch model and the input resolution while ensuring that the total computational cost is close. Specifically, when using the same input image resolution, the combination of models of different sizes does not bring significant improvements to detection performance. Correspondingly, when the three branches use the same model (e.g. BBB), the input image resolution is adjusted to the pyramid structure. The performance of the final model is slightly improved on $\rm AP^{b}$ (44.5% vs. 43.8%), but the $\rm AP^{m}$ drops significantly (38.7% vs 39.9%) due to the reduction of the maximum resolution. The former demonstrates the importance of the image pyramid, and the latter further demonstrates the need for the image pyramid to maintain a larger image scale range, which is especially essential for instance segmentation. Drawing on experience, parameter-inverted image networks are an efficient design method that can meet the above requirements, especially when compared to its opposite configuration parameter-direct image pyramid, i.e. TSB with 448/672/896 resolution (46.6% vs. 42.6%). As shown in Table 9, with less computation than the baseline, the model can support image inputs in the maximum range from 672 to 1,568, and the performance is significantly improved.
@@
Branch merging. Experiments in Table 8 prove that branch merging of all branches yields the best performance by providing multi-scale semantically rich features, compared to only using feature maps from single or partial branches.
Attention type. The core of information interaction between branches is cross-attention. We adopt PIIP-TSB with resolution 1120/896/448 as the basic model and investigate two different attention mechanisms. As shown in Table 10, deformable attention [51] with linear complexity can significantly improve the performance of the model without substantially increasing the computational cost. We end up using deformable attention as the default configuration. Notably, it can be replaced by other more advanced attention mechanisms in the future to further boost performance.
@@
Number of interactions. As shown in Table 10, no matter which attention mechanism is used, the increase in the number of interactions will improve the performance of the model to varying degrees. Since it also increases the computational cost, we further explore the cost-effectiveness of different numbers of interactions. We conduct experiments with different resolution combinations on models with different numbers of interactions, and the scatter plot of all results is shown in Fig. 5(b). It can be seen that when the number of interactions is small (less than 2), the growth trend of model performance with the increase in computational cost is relatively slow. We attribute this to too few interactions and insufficient information complementation between branches. Therefore, we use 12 interactions by default. Note that as the model size increases (e.g. more layers), the number of interactions can also increase accordingly.
@@
Interaction direction between branches. We compare five different interaction directions in Table 11. Considering both the computational cost and performance, we finally choose the fourth method, i.e. bidirectional connections of adjacent branches, as the default choice. As can be seen from the scatter plot in Table 11, all the interaction directions achieve a satisfactory performance-computation balance, validating their ability to improve communication between branches.
@@ 5 Conclusion
Appendix A Detailed Training Settings for Image Classification
@@
Appendix B Full Detection Results
@@
Appendix C Preliminary Attempt on From-scratch Pre-training

### v2

# Xizhou Zhu2,1*,  Xue Yang1*,  Zhaokai Wang3,1*,  Hao Li4,1 Wenhan Dou2,5,  Junqi Ge2,5,  Lewei Lu5,  Yu Qiao1,  Jifeng Dai2,1†

@@ Abstract
00footnotetext: Equal contribution. Corresponding author: Jifeng Dai <daijifeng@tsinghua.edu.cn>.
@@ 4.1 Implementation Details
For comparison with Base-size models, we use pre-trained ViT-T/S/B as the branches to construct three-branch PIIP network, namely PIIP-TSB. Similarly, ViT-S/B/L are used to construct PIIP-SBL to match the computation of Large-size models. We also construct four-branch PIIP-TSBL with ViT-T/S/B/L. We set the number of interactions (each with 2 interaction units as shown in Fig. 2) $N$ to 12, i.e. after every layer for ViT-T/S/B or after every two layers for ViT-L. We construct multiple variants of three-branch and four-branch models with different resolution configurations. For combinations with an inconsistent number of layers, we will use a larger learning rate decay for the backbone with fewer layers. For example, for ViT-S/B (12 layers) and ViT-L (24 layers), the learning rate decay for ViT-S/B is set to be twice that of ViT-L (24/12=2).
@@
We use the FLOPs calculation script from MMDetection [6], with our modifications to accurately calculate FLOPs of modules like self-attention and deformable attention. The script is released along with the training code. We have also manually verified the calculations using formulas, and the results are consistent with those produced by the script.
@@ 4.2 Object Detection and Instance Segmentation
Table 2: Object detection and instance segmentation performance on COCO val2017. ‘MS’ means using AutoAugment [10] for multi-scale training. Large-size models use ViT weights trained on ImageNet-21K. The ViTDet-B and ViTDet-L results (and other entries) are cited from ViT-Adapter [7]. PIIP-SBL with Mask R-CNN uses higher resolutions than those in Tab. 1, as reported in Tab. 14. For PIIP-TSB with Mask R-CNN, higher resolutions (1568/896/672 -> 1792/1344/672) and a larger window size (14 -> 28) are used, compared with the results in the Tab. 1.
@@
DINO + MS schedule
PIIP-SBL-3$\times$ (ours) 57.9 76.9 63.3 - - -
PIIP-H6B-1$\times$ (ours) 60.0 79.0 65.4 - - -
@@
Effectiveness of Parameter-Inverted Image Pyramid. To demonstrate the performance and computational advantages of the Parameter-Inverted Image Pyramid (PIIP) Networks, we perform validation on two baseline models ViTDet-B and ViTDet-L [23] in Tab. 1. Taking the three-branch structure as an example, while maintaining similar performance with ViTDet-B, our PIIP-TSB reduces the computational cost by 47.5% (243G vs. 463G) and 38.0% (287G vs. 463G) in object detection and instance segmentation tasks respectively. Similarly, compared with ViTDet-L, our PIIP-SBL reduces the computational cost by about 52.9% (727G vs. 1,542G) and 35.0% (1,002G vs. 1,542G) in the above two tasks respectively. On the other hand, with similar computational cost as the baseline, PIIP-TSB and PIIP-SBL improve the object detection performance by 2.8% and 2.6%, respectively, and instance segmentation by 1.5% and 1.2%, compared to ViTDet-B and ViTDet-L. To better illustrate the above conclusion, we depict the trend between the computational cost and performance of different PIIP model combinations by adjusting the input resolution, as shown in Fig. 4. Furthermore, when we use the four-branch structures, the curve in the figure is slightly better than that of the three-branch structure.
@@
Results with Base-size and Large-size models. As shown in Tab. 2, combined with Mask R-CNN, PIIP achieves higher performance than ViT-Adapter by a considerable margin, about 0.9% and 1.2% on $\rm AP^{b}$. With a more powerful detector Cascade R-CNN and stronger training schedule (3$\times$ + MS), PIIP-TSB and PIIP-SBL achieve competitive performance of 53.1% and 54.5% $\rm AP^{b}$, respectively. Finally, we achieve 57.9% $\rm AP^{b}$ with the DINO [61] detector.
Results with InternViT-6B. We further examine PIIP on an extremely large vision foundation model InternViT-6B [8], and finally achieve 55.7% $\rm AP^{b}$ by using Mask R-CNN 1$\times$ training schedule. In addition, as can be seen from Tab. 3, our PIIP can save nearly 43% of the computation and achieve better performance than single-branch InternViT-6B by 1.9% on $\rm AP^{b}$ and 0.9% on $\rm AP^{m}$.
Results with different pre-training methods. We initialize PIIP-SBL with pre-trained ViT S/B/L weights from AugReg [43], Uni-Perceiver [66], MAE [18], DeiT III [46], DINOv2 [38] and BEiTv2 [39]. As shown in Tab. 6, the BEiTv2-initialized model achieves the best performance.
@@ 4.3 Semantic Segmentation
Results with Base-size and Large-size models. In Tab. 6, PIIP can achieve better performance with fewer computations. Moreover, PIIP-TSB in Tab. 6 attains 51.6% mIoU with UperNet, exceeding InternImage-B [50] by 1.4%. Similarly, PIIP-SBL yields the results of 54.3% mIoU, which is outstanding compared to counterparts like ConvNeXt-XL [31] and InternImage-L [50].
@@
Table 8: Ablation on Branch Merging on COCO val2017. We use PIIP-TSB 1568/896/672.
@@ 4.4 Image Classification
Results. As shown in Tab. 8, when compared with the DeiT baseline, our PIIP-SBL reduces the computational cost by 36.7% (39.0G vs. 61.6G) while maintaining the performance. When using a similar computational cost as the baseline models, PIIP-TSB and PIIP-SBL improve the top-1 accuracy by 0.3% and 0.7%, respectively.
Table 9: Ablation on image pyramid and parameter-inverted design. ‘PI’, ‘IP’ and ‘Inter.’ represent parameter-inverted, image pyramid and interactions. ‘MS’ means multi-scale training, following [10].
@@ 4.5 Ablation Study
Superiority of parameter-inverted image networks. We evaluate the effectiveness of the image pyramid and parameter-inverted design by comparing our method with other methods, e.g. designs in Fig. 1. First of all, a single-branch with multi-scale training is the simplest image pyramid practice, as shown in Tab. 9. Compared with the baseline model, its performance improvement is limited (44.8% vs. 43.8%). Secondly, we conduct experiments by controlling the scale of the branch model and the input resolution while ensuring that the total computational cost is close. Specifically, when using the same input image resolution, the combination of models of different sizes does not bring significant improvements to detection performance. Correspondingly, when the three branches use the same model (e.g. BBB), the input image resolution is adjusted to the pyramid structure. The performance of the final model is slightly improved on $\rm AP^{b}$ (44.5% vs. 43.8%), but the $\rm AP^{m}$ drops significantly (38.7% vs 39.9%) due to the reduction of the maximum resolution. The former demonstrates the importance of the image pyramid, and the latter further demonstrates the need for the image pyramid to maintain a larger image scale range, which is especially essential for instance segmentation. Drawing on experience, parameter-inverted image networks are an efficient design method that can meet the above requirements, especially when compared to its opposite configuration parameter-direct image pyramid, i.e. TSB with 448/672/896 resolution (46.6% vs. 42.6%). As shown in Tab. 9, with less computation than the baseline, the model can support image inputs in the maximum range from 672 to 1,568, and the performance is significantly improved.
Table 10: Baseline with higher resolution.
Model Resolution #Param #FLOPs $\rm AP^{b}$
ViTDet-L 1024 308M 1542G 46.8
ViTDet-L 1792 308M 6458G 48.3
PIIP-TSBL 1792/1568/1120/448 512M 1535G 49.6
Baseline with higher resolution. To evaluate the impact of using higher resolutions, we add another baseline, ViTDet-L with 1792 resolution (matching the largest resolution of PIIP-TSBL 1792/1568/1120/448), as shown in the second row of the Tab. 10. The first and third rows are from Table 1. We observed that while ViTDet-L with 1792 resolution achieves better performance compared to the 1024 resolution, its FLOPs are approximately 4 times larger. Compared with PIIP-TSBL, ViTDet-L 1792 has a lower box AP (-1.3%) and 4 times larger FLOPs. This experiment well explains that the performance improvement does not entirely come from larger image resolution. The structure of PIIP also plays an important contribution in the calculation amount and performance improvement.
@@
Branch merging. Experiments in Tab. 8 prove that branch merging of all branches yields the best performance by providing multi-scale semantically rich features, compared to only using feature maps from single or partial branches.
Attention type. The core of information interaction between branches is cross-attention. We adopt PIIP-TSB with resolution 1120/896/448 as the basic model and investigate two different attention mechanisms. As shown in Tab. 11, deformable attention [55] with linear complexity can significantly improve the performance of the model without substantially increasing the computational cost. We end up using deformable attention as the default configuration. Notably, it can be replaced by other more advanced attention mechanisms in the future to further boost performance.
@@
Number of interactions. As shown in Tab. 11, no matter which attention mechanism is used, the increase in the number of interactions will improve the performance of the model to varying degrees. Since it also increases the computational cost, we further explore the cost-effectiveness of different numbers of interactions. We conduct experiments with different resolution combinations on models with different numbers of interactions, and the scatter plot of all results is shown in Fig. 5(b). It can be seen that when the number of interactions is small (less than 2), the growth trend of model performance with the increase in computational cost is relatively slow. We attribute this to too few interactions and insufficient information complementation between branches. Therefore, we use 12 interactions by default. Note that as the model size increases (e.g. more layers), the number of interactions can also increase accordingly.
@@
Interaction direction between branches. We compare five different interaction directions in Tab. 12. Considering both the computational cost and performance, we finally choose the fourth method, i.e. bidirectional connections of adjacent branches, as the default choice. As can be seen from the scatter plot in Tab. 12, all the interaction directions achieve a satisfactory performance-computation balance, validating their ability to improve communication between branches.
@@ 5 Conclusion
Limitations. While our method manages to save computation, its memory consumption is higher than single-branch models due to the increase of parameter count. Our current method only focuses on ViT-based models. PIIP with hierarchical networks (e.g. CNN) or heterogeneous structures (e.g. CNN for some branches and ViT for other branches) remain unexplored for future work.
Acknowledgement
This work is supported by the National Key R&D Program of China (NO. 2022ZD0161300, NO. 2022ZD0160100), by the National Natural Science Foundation of China (62376134).
Appendix A Appendix
A.1 Detailed Training Settings for Image Classification
@@
A.2 Full Detection Results
@@
A.3 Preliminary Attempt on From-scratch Pre-training
@@
A.4 Broader Impacts
Our method helps to save computational overheads of large-scale vision foundation models such as InternViT-6B, therefore reducing energy consumption. This may bring positive impacts on carbon emissions reduction and contribute to environmental sustainability. However, energy consumption of large models still needs to be treated with caution.
A.5 Licenses of Datasets
ImageNet-1k [11] is subject to the ImageNet terms of use [47].
COCO [28] is subject to the Flickr terms of use [15].
ADE20K [64] is subject to the ADE20K terms of use [36].